OpenStudy (anonymous):

\sqrt[4]{6}+\sqrt[2]{6}-\sqrt{6} answer please. I know this site isnt meant to cheat but please. Help a guy with his homework

OpenStudy (unklerhaukus):

\[\sqrt[4]{6}+\sqrt[2]{6}-\sqrt{6}\]

OpenStudy (unklerhaukus):

\(\sqrt[2]{}\) is the same as \(\sqrt{}\)

OpenStudy (anonymous):

can you simplify it?

OpenStudy (anonymous):

\[\sqrt[4]{6}+\sqrt[2]{6}-\sqrt{6}\] simplify fully

OpenStudy (mathstudent55):

The part in red shows like terms that can be combined together. \(\large \sqrt[4]{6}+\color{red}{\sqrt[2]{6}-\sqrt{6}}\)